Hook Industries's capital structure consists solely of debt and common equity. It can issue debt at rd = 10%, and its common stock currently pays a $1.50 dividend per share (D0 = $1.50). The stock's price is currently $32.50, its dividend is expected to grow at a constant rate of 9% per year, its tax rate is 35%, and its WACC is 13.55%. What percentage of the company's capital structure consists of debt? Do not round intermediate calculations. Round your answer to two decimal places. %

I am having problems with this question please help
Answer #1

Cost of equity=(D1/Current price)+Growth rate

=(1.5*1.09)/32.5+0.09

=14.03076923%

Cost of debt after tax=10*(1-tax rate)

=10(1-0.35)=6.5%

Let debt be $x

Equity be $y

Total=$(x+y)

WACC=Respective costs*Respective weight

13.55=(x/(x+y)*6.5)+(y/(x+y)*14.03076923)

13.55*(x+y)=6.5x+14.03076923y

13.55x+13.55x=6.5x+14.03076923y

x=(14.03076923-13.55)y/(13.55-6.5)

=0.068194217y

Total=x+y

=1.068194217y

Hence weight of debt=0.068194217y/1.068194217y

=6.38%(Approx).